OpenStudy (gabbyalicorn):

Please help! *will medal* Thanks! solve for x. drawing pic give me a sec lower left is 42º lower right is (5x+4)º upper right is (7x+6)º

OpenStudy (gabbyalicorn):



OpenStudy (candy13106):

they r vertical angle so they have to equal to each other

OpenStudy (anonymous):

The angles in a triangle add up to 180, so we can build an equation to solve: \(180=42+(5x+4)+(7x+6)\) \(180=52+12x\) Can you solve this for x?

OpenStudy (gabbyalicorn):

for the first one it's 32/3? and the second one is 32/3

OpenStudy (gabbyalicorn):

so x = 32/3 ? @tom982

OpenStudy (anonymous):

Yes, \(x=\frac{32}{3}\). Good job.

OpenStudy (gabbyalicorn):

cool :D

OpenStudy (gabbyalicorn):

ty so much!!!!